# Training a Utility-based Retriever Through Shared Context Attribution for Retrieval-Augmented Language Models

### v1

@@ Abstract
Retrieval-Augmented Language Models boost task performance, owing to the retriever that provides external knowledge. Although crucial, the retriever primarily focuses on semantics relevance, which may not always be effective for generation. Thus, utility-based retrieval has emerged as a promising topic, prioritizing passages that provides valid benefits for downstream tasks. However, due to insufficient understanding, capturing passage utility accurately remains unexplored. This work proposes SCARLet, a framework for training utility-based retrievers in RALMs, which incorporates two key factors, multi-task generalization and inter-passage interaction. First, SCARLet constructs shared context on which training data for various tasks is synthesized. This mitigates semantic bias from context differences, allowing retrievers to focus on learning task-specific utility for better task generalization. Next, SCARLet uses a perturbation-based attribution method to estimate passage-level utility for shared context, which reflects interactions between passages and provides more accurate feedback. We evaluate our approach on ten datasets across various tasks, both in-domain and out-of-domain, showing that retrievers trained by SCARLet consistently improve the overall performance of RALMs.
@@
Yilong Xu1,2,3 Jinhua Gao1,2  Xiaoming Yu1,2 Yuanhai Xue1,2  Baolong Bi1,2,3 Huawei Shen1,2,3Xueqi Cheng1,2,3 1State Key Lab of AI Safety, Institute of Computing Technology, CAS 2Key Lab of AI Safety, Chinese Academy of Sciences 3University of Chinese Academy of Sciences {xuyilong23s, gaojinhua}@ict.ac.cn
@@ 1 Introduction
Early RALMs adopt relevance-based retrievers, including both sparse (Robertson and Zaragoza, 2009) and dense (Karpukhin et al., 2020) models. However, these retrievers are primarily biased toward semantic relevance, failing to capture the utility of retrieved passages (Wu et al., 2024). Utility refers to the valid gain that a passage contributes to the downstream task generation (Zhang et al., 2024), and is considered a better indicator suited for RALMs. Some recent works explore to optimize retrievers by constructing feedback from generators (Shi et al., 2023; Yu et al., 2023; Wei et al., 2024), achieving promising results. Nonetheless, how to better align retrievers with generators to capture the utility of retrieved passages remains an open yet challenging problem.
Different from relevance which is mainly determined by the query and the passage, the utility needs more comprehensive measurement. In this paper, we propose the following two vital yet overlooked factors for utility modeling in RALMs:
@@
RALMs need to accommodate various downstream tasks, where the utility of a passage can vary accordingly. Existing methods typically optimize retrievers in the pooled setup, i.e., mixing data from different datasets for training, to learn task-specific retrieval criteria (Lin et al., 2024; Zamani and Bendersky, 2024). However, due to contextual difference, these exist huge semantic gaps across datasets, resulting in insufficient training where the semantic relevance might outweigh the utility features, which can affect generalization, especially for retrievers with weaker linguistic capabilities (Liu et al., 2024).
@@
In some complex tasks, the utility of a certain passage cannot be solely determined by itself. For example, when handling multi-hop question-answering tasks, the model should rely on preceding and even succeeding contexts on the reasoning chain to judge the utility of a certain passage. However, the utility signals constructed in previous works either fail to capture the fine-grained feedback at the passage level (Zamani and Bendersky, 2024; Sohn et al., 2024) or evaluate each passage independently (Yu et al., 2023; Shi et al., 2023).
To address the limitations in the above two aspects and improve utility-based retrieval, we propose a novel framework named SCARLet, which represents shared context attribution supervised training for utility-based retrievers.
Specifically, SCARLet first introduces a data synthesis pipeline. Contrary to previous approaches, this pipeline first constructs a shared context, and subsequently synthesizes training data for various downstream tasks derived from the shared context. This method mitigates the semantic interference by achieving single-variable control, thereby enabling the retriever to focus on learning task-specific utility. During the downstream generation stage, SCARLet attributes utility to the shared context using a passage-level perturbation-based technique, which measures each passage’s contribution by removing or retaining passages from the context, then observing the fluctuations in the generator’s output. This approach effectively captures the synergy between passages, thereby accurately reflecting their utility. Leveraging the shared context with utility signals, SCARLet uses one-dimensional clustering to sample positive and negative passages, tuning the retriever to shift the retrieval criterion toward utility.
We evaluate the improvement that SCARLet brings to the overall performance of RALMs. We compare several retrievers and previous approaches that optimize retrievers based on feedback from generation. Our experiments select ten datasets, both in-domain and out-of-domain, covering eight distinct tasks. The results show that generators augmented with retrievers trained by SCARLet consistently achieves optimal or suboptimal downstream performance across all datasets. Additionally, we conduct retrieval evaluation and case study to further demonstrate the effectiveness of SCARLet.
@@
  • We argue that utility-based retrieval needs to consider the following two factors: multi-task generalization and inter-passage interaction.
  • We propose SCARLet, a novel framework that optimizes the retriever through shared context synthesis and utility attribution.
  • We conduct experiment across various tasks, demonstrating that our proposed SCARLet improves the overall performance of RALMs.
@@ 2 Related Work
In RALMs, by the definition of utility, we review related works that optimizing retrievers from the feedback of generation, which are essentially explorations of retrieval based on downstream utility. Salemi and Zamani (2024) propose supervision based on downstream task metrics, but fail to provide passage-level utility feedback. Shi et al. (2023); Yu et al. (2023) assess utility of each passage using generator outputs, but they ignore the the interactions between passages. Sohn et al. (2024); Wei et al. (2024) employ the generator’s self-reflection to evaluate utility, which may bring hallucinations as the language models can be dishonest  (Madsen et al., 2024). Asai et al. (2023); Glass et al. (2022) notice the multi-task nature of the retrieval stage, but fail to account for the training biases introduced by separate datasets.
Therefore, our SCARLet framework comprehensively considers the above task-level and passage-level limitations, offering a pipeline of shared context synthesis and attribution to effectively train utility-based retrieval in RALMs.
@@ 3.1 Definitions
In a typical RALM system, there is a retriever and a generator. During the retrieval stage, we employ a dense retriever, assuming that it is based on an encoder $\mathbf{Enc}$ with parameters $\phi$. Additionally, the retriever interacts with an external corpus $\mathcal{C}$.
In a downstream task $T$, for an input query $x$, we calculate the dot product of the embeddings of $x$ and each passage $d$ in $\mathcal{C}$, as follows:
$\mathrm{score}\left(x,d\right)=\mathbf{Enc}_{\phi}\left(x\right)\cdot\mathbf{Enc}_{\phi}\left(d\right),d\in\mathcal{C}.$ (1)
The top-$k$ passages with the highest scores are selected and added to the context, denoted as $D=\left[d_{1},...,d_{k}\right]$. Given that we consider multiple downstream tasks, the form of retrieval query is modified to $x=I+x$, where $I$ represents the instruction description of task $T$. In the generation stage, a language model $\mathrm{LM}$ with parameters $\theta$ typically serves as the generator. The context $D$ is used to enhance the generation, ultimately producing the predicted output $\mathbf{\hat{y}}$, as shown below:
$\mathbf{\hat{y}}=\mathrm{LM}_{\theta}\left(x\oplus D\right),$ (2)
@@ 3.2 Overview of SCARLet
Our proposed SCARLet mainly consists of three parts, including the synthetic pipeline for shared context and data (§3.3), the method for utility attribution (§3.4), as well as the data sampling and the training objective for tuning the retriever (§3.5). Please see Figure 1 for a flowchart of the method.
Shared context means that in RALMs, during training stage, data from different tasks share the same context which includes multiple passages. Previous methods utilize different training datasets with varying corpora (Lin et al., 2024; Zamani and Bendersky, 2024). However, context differences can interfere with retrievers in learning task-specific utility, especially for models with weaker linguistic capabilities. Furthermore, existing strategies for mitigating semantic bias typically construct noisy context, which first retrieves useful passages from training data and adds noisy passages manually or semi-manually, resulting in low recall rates for useful passages and high labor costs. To tackle the above challenges, our proposed SCARLet adopts a reverse strategy, first constructing shared context to narrow the semantic gap, and then synthesizing task-specific data based on this context. Sharing context across tasks can highlight utility feature differences, making it easier to learn. And LLM-driven data synthesis has shown to be a promising way (Long et al., 2024; Kim and Baek, 2025), which can effectively reduce labor costs.
Attribution refers to local explanation techniques that provide evidence for the model’s output (Li et al., 2023). And contributive attribution measures the contribution of the source data to the generation (Worledge et al., 2023), with higher scores indicating a greater positive impact. This aligns well with the definition of utility we are studying, when using the retrieved context as the source data for downstream generation in RALMs. Previous research on optimizing retrievers from downstream generation, either fails to construct passage-level feedback or only consider the individual impact of each passage, overlooking the synergy effects between passages. Therefore, taking the shared context as the source data, SCARLet uses a passage-level perturbation-based attribution approach, where variations in generation caused by perturbations can reflect interactions between passages.
@@ 3.3 Shared Context Synthesis
Specifically, we begin by collecting seed data from existing datasets, including task instructions, inputs, and ground truth. In line with the motivation behind shared context, passages within this context need to be closely related to facilitate the synthesis of high-quality data. Therefore, we employ an approach based on associated entities, which extract entities from the seed data, search their adjacent entities by querying Wikidata [^1] https://www.wikidata.org, and merge them to obtain a related entity list. Then we use this list to retrieve relevant passages from the Wikipedia corpus, which then serve as the shared context $D_{\mathrm{shared}}$. Subsequently, we instruct the synthesizer model $\mathsf{S}$ to generate new training data, using $D_{\mathrm{shared}}$ as the information source. We also input the task pool $\mathcal{T}$, which is linked to various downstream tasks, such as multi-hop QA, long-form QA, and fact checking. The process is formalized as follows:
@@
To improve the quality of synthetic data, the task pool $\mathcal{T}$ not only provides the task instructions but also offers example data. To further improve robustness, following Fang et al. (2024); Zhang et al. (2024), we also introduce synthetic noise into the shared context by instructing the synthesizer to generate semantically relevant but useless passages. In addition, we incorporate a data filtering step that instructs the synthesizer to check for and eliminate new data containing faults. For more details, please refer to Appendix A. We also provide an example of shared context in Appendix F.
@@
3.4 Passage-level Utility Attribution
Specifically, by definition, the context $D$ recalled by the upstream retrieval consists of $k$ passages. The perturbation-based method modifies the context using a perturbation vector $\mathbf{v}\in\left\{0,1\right\}^{k}$, where 0 and 1 indicate whether the corresponding passage is removed or included, respectively. However, running all generations of $2^{k}$ possible perturbation vectors can result in significant computational overhead. Inspired by the method of Local Interpretable Model-agnostic Explanations (LIME; Ribeiro et al., 2016; Mardaoui and Garreau, 2021), we first sample $n$ perturbation vectors randomly and then fit a surrogate model for predicting the utility score, as shown below:
where we adopt the ridge regression (Hilt and Seegrist, 1977) as our surrogate model, $\bm{\alpha}$ represents the parameters to be fitted, $\lambda$ is a hyperparameter for regularization, and $z_{i}$ is the observed value under $\mathbf{v}_{i}$. More detailed, $\bm{\alpha}^{(i)}$ denotes the utility score of passage $d_{i}$, $\bm{\alpha}^{(0)}$ represents the intercept term. And $z_{i}$, which quantifies the fluctuation caused by $\mathbf{v}_{i}$, is calculated using the logit values of the tokens in the ground truth $\mathbf{y}$ at each time step, as shown below:
To evaluate the effectiveness of the above utility attribution method, we conduct a preliminary experiment on the GTI benchmark (Zhang et al., 2024), which includes three datasets: HotpotQA  (Yang et al., 2018), Natural Questions  (NQ; Kwiatkowski et al., 2019), and MSMARCO-QA  (Bajaj et al., 2018). Each test data comprises input, ground truth, and ten passages including correct passages and other noise passages. We use the utility score to rank the passages. The results, measured using the nDCG, demonstrate that our method shows a high accuracy in reflecting passage utility, as shown in Figure 2. We also compare our method to other attribution approaches, and our method outperforms them by over 20%. For further details of the experiment, please refer to Appendix B.
@@ 4.1 Datasets and Evaluation
The datasets we select are categorized into in-domain and out-of-domain. In-domain datasets provide training splits for constructing seed data in SCARLet, while out-of-domain datasets are collected for generalization testing since factors such as task, data domain, and corpus are critical external influences on RALM performance. We collect seven datasets from KILT (Petroni et al., 2021), and three from BeIR (Thakur et al., 2021), as detailed in Table 1. All KILT datasets utilize Wikipedia as the corpus $\mathcal{C}$, specifically the dump dated 2019-08-01222http://dl.fbaipublicfiles.com/BLINK/enwiki-pages-articles.xml.bz2. And following Wang et al. (2019), we split the original articles into segments with a maximum length of 100 words, resulting in a total of 28,773,800 passages. The datasets from BeIR utilize their own dedicated corpus. For retrieval, we follow the closed setup(Asai et al., 2023), where RALMs only retrieve from the corpus of the current dataset. For the test data, we randomly sample 1,000 data from the test split of each dataset.
For evaluation metrics, since the primary focus of this study is on utility-based retrieval, which pursues overall improvements of RALMs, we mainly assess the performance of downstream tasks. For WoW, we use F1. For ELI5 and FiQA, we use ROUGE-L. For other datasets, we use accuracy.

### v2

@@ Abstract
Retrieval-Augmented Language Models boost task performance, owing to the retriever that provides external knowledge. Although crucial, the retriever primarily focuses on semantics relevance, which may not always be effective for generation. Thus, utility-based retrieval has emerged as a promising topic, prioritizing passages that provide valid benefits for downstream tasks. However, due to insufficient understanding, capturing passage utility accurately remains unexplored. This work proposes SCARLet, a framework for training utility-based retrievers in RALMs, which incorporates two key factors, multi-task generalization and inter-passage interaction. First, SCARLet constructs shared context on which training data for various tasks is synthesized. This mitigates semantic bias from context differences, allowing retrievers to focus on learning task-specific utility and generalize across tasks. Next, SCARLet uses a perturbation-based attribution method to estimate passage-level utility for shared context, which reflects interactions between passages and provides more accurate feedback. We evaluate our approach on ten datasets across various tasks, both in-domain and out-of-domain, showing that retrievers trained by SCARLet consistently improve the overall performance of RALMs.
Resources: github.com/ylXuu/SCARLet
@@
Yilong Xu1,2,3  Jinhua Gao1,2thanks: Corresponding author.   Xiaoming Yu1,2  Yuanhai Xue1,2   Baolong Bi1,2,3 Huawei Shen1,2,3Xueqi Cheng1,2,3 1State Key Lab of AI Safety, Institute of Computing Technology, CAS 2Key Lab of AI Safety, Chinese Academy of Sciences 3University of Chinese Academy of Sciences {xuyilong23s, gaojinhua}@ict.ac.cn
@@ 1 Introduction
Early RALMs adopt relevance-based retrievers, including both sparse (Robertson and Zaragoza, 2009) and dense (Karpukhin et al., 2020) models. However, these retrievers are primarily biased toward semantic relevance (Wu et al., 2024), failing to consider the passage utility and leading to misalignment in RALMs. The utility, measuring the valid gain that a passage contributes to the downstream generation (Zhang et al., 2024), can bridge the gap between the retriever and the generator. Some recent works have proposed to optimize retrievers by constructing feedback from generators (Shi et al., 2023; Yu et al., 2023; Wei et al., 2024), achieving promising results. Nonetheless, how to align retrievers to better capture utility remains an open yet challenging problem.
Different from relevance, which is mainly determined by the query and the passage (Xia et al., 2015; Guo et al., 2019), utility needs a more comprehensive measurement. In this paper, we propose the following two vital yet overlooked factors for utility modeling in RALMs:
@@
RALMs need to accommodate various downstream tasks, where the utility of a passage can vary accordingly. Existing methods typically optimize retrievers using the pooling strategy, i.e., mixing data from different tasks for training, to learn task-specific retrieval criteria (Lin et al., 2024; Zamani and Bendersky, 2024). However, since pooled samples from different tasks typically have different contexts, the trained retrievers might tend to capture semantic relevance signals instead of utility features. Such unexpected preference will downgrade the retrievers’ generalization ability, especially for those with weaker linguistic capabilities (Liu et al., 2024).
@@
In some complex tasks, the utility of a certain passage cannot be solely determined by itself. For example, when handling multi-hop question-answering tasks, the model should rely on preceding and even succeeding contexts in the reasoning chain to judge a passage’s utility. However, the utility signals constructed in previous works either fail to provide passage-level feedback (Zamani and Bendersky, 2024; Sohn et al., 2024) or evaluate each passage independently (Yu et al., 2023; Shi et al., 2023), leading to imprecise utility measurements.
In this paper, we propose a novel framework to train utility-based retrievers for RALMs, named SCARLet, representing shared context attribution supervised training for utility-based retrievers.
Specifically, SCARLet first introduces a training data synthesis pipeline. Contrary to the previous pooling strategy that mixes training data with different contexts, our pipeline first constructs a shared context, and subsequently synthesizes training data for various downstream tasks derived from the shared context. This method mitigates the semantic interference by achieving single-variable control, and enables the retriever to focus on learning task-specific utility. To better assess the utility of certain passages, SCARLet employs a passage-level perturbation-based technique, which randomly removes some passages from the context and measures the fluctuations in the generated output. Such a design can effectively capture the synergy between passages, thereby accurately reflecting their utility. Finally, SCARLet collects positive and negative samples based on the utility scores and trains the retriever in a contrastive way.
We conduct extensive experiments to evaluate the performance gain brought by SCARLet. Our experiments adopt ten datasets, covering eight distinct tasks that are frequently used for RALMs evaluation. The results show that RALMs equipped with retrievers trained by SCARLet, consistently achieve optimal or suboptimal downstream performance across all datasets. Moreover, further analysis and case studies demonstrate that SCARLet can better capture utility signals.
@@
  • We argue that utility should be preferred in RALMs and propose two critical factors for training utility-based retrievers.
  • We propose SCARLet, a novel framework to train utility-based retrievers through shared context synthesis and utility attribution.
  • We conduct extensive experiments across various tasks, demonstrating that our proposed SCARLet can improve the overall performance of RALMs.
@@ 2 Related Work
In RALMs, early exploration of retrieval utility focuses on capturing the downstream feedback of generators. Salemi and Zamani (2024) propose supervision based on downstream task metrics, but fail to provide passage-level utility feedback. Shi et al. (2023); Yu et al. (2023) assess utility of each passage using generator outputs, but they ignore the interactions between passages. Sohn et al. (2024); Wei et al. (2024) employ the generator’s self-reflection to evaluate utility, which may bring hallucinations as the language models can be dishonest  (Madsen et al., 2024). Asai et al. (2023); Glass et al. (2022) notice the multi-task nature of the retrieval stage, but fail to account for the training biases introduced by contextual differences in the pooling strategy.
Therefore, our proposed SCARLet framework comprehensively considers the above issues of multi-task generalization and utility assessment, offering a novel pipeline with shared context synthesis and utility attribution to effectively train utility-based retrievers in RALMs.
@@ 3.1 Definitions
A typical RALM system consists of a retriever and a generator. During the retrieval stage, we employ a dense retriever based on an encoder $\mathbf{Enc}$ with parameters $\phi$. And the retriever interacts with an external corpus $\mathcal{C}$. For a query $q$, we calculate the dot product of the embeddings of $q$ and each passage $d$ in $\mathcal{C}$, as follows:
$\mathrm{score}\left(q,d\right)=\mathbf{Enc}_{\phi}\left(q\right)\cdot\mathbf{Enc}_{\phi}\left(d\right),d\in\mathcal{C}.$ (1)
The top-$k$ passages with the highest scores are selected and added to the context, denoted as $D=\left[d_{1},\dots,d_{k}\right]$. Note that RALMs need to accommodate various downstream tasks, for a task $T$ and an input $x$ from its dataset, we define the query format as $q=I\oplus x$, where $I$ denotes the instruction description of task $T$.
In the generation stage, a language model $\mathrm{LM}$ with parameters $\theta$ serves as the generator. The context $D$ is used to enhance generation, ultimately producing the predicted output $\mathbf{\hat{y}}$, as shown below:
$\mathbf{\hat{y}}=\mathrm{LM}_{\theta}\left(I\oplus x\oplus D\right),$ (2)
@@ 3.2 Overview of SCARLet
The overall architecture of our proposed SCARLet is shown in Figure 1, including shared context synthesis and training data construction (§3.3), utility attribution modeling (§3.4), as well as data sampling and retriever tuning (§3.5).
Shared context refers to the common context for data of different tasks in the training stage, which is then used to enhance downstream generation. Previous studies employ the pooling strategy (Lin et al., 2024; Zamani and Bendersky, 2024), where each instance has a distinct context for training. Learning task-specific features to improve multi-task generalization of utility-based retrieval might be disturbed by the semantically relevant noise introduced by differences in context, leading to unexpected preference, particularly in retrievers with weaker linguistic capabilities. To tackle the above challenges, our proposed SCARLet adopts a reverse strategy, first constructing shared context to narrow the semantic gap, and then synthesizing task-specific data based on this context. Sharing context across tasks can highlight utility feature differences, making it easier to learn. Moreover, LLM-driven data synthesis has been shown to be a promising way (Long et al., 2024; Kim and Baek, 2025), which can effectively reduce labor costs.
Utility attribution modeling refers to local explanation techniques to build utility signals from the downstream generation. More specifically, we adopt the contributive attribution model, which measures how the input context contributes to the model’s output and aligns well with the definition of utility in RALMs. Previous research on optimizing retrievers from downstream generation, either fails to construct passage-level feedback or only considers the individual impact of each passage, overlooking the synergy effects between passages. Therefore, taking the shared context as the source data, SCARLet uses a passage-level perturbation-based utility attribution approach, where fluctuations in generation caused by perturbations can reflect interactions between passages and then be quantified as utility scores.
@@ 3.3 Shared Context Synthesis
Specifically, we first define a task pool $\mathcal{T}$, which is linked to various downstream tasks and their datasets, such as multi-hop QA, long-form QA, and fact checking. We begin by collecting seed data from datasets of $\mathcal{T}$, including task instructions, inputs, and ground truth. In line with the motivation behind shared context, passages within this context need to be closely related to facilitate the synthesis of high-quality data. Therefore, we employ an approach based on associated entities, which extracts entities from the seed data, searches their adjacent entities by querying Wikidata [^1] https://www.wikidata.org, and merges them to obtain a related entity list. We then use this list to retrieve relevant passages from the Wikipedia corpus, and treat the recalled passages as the shared context $D_{\mathrm{shared}}$. Subsequently, we instruct the synthesizer model $\mathsf{S}$ to generate new training data, using $D_{\mathrm{shared}}$ as the information source and task information (including instructions and examples) from $\mathcal{T}$. The process is formalized as follows:
@@
To improve the quality of synthetic data, the task pool $\mathcal{T}$ not only provides the task instructions but also offers example data. To further improve robustness, following Fang et al. (2024); Zhang et al. (2024), we also introduce synthetic noise into the shared context by instructing the synthesizer to generate semantically relevant but useless passages. In addition, we incorporate a data filtering step that instructs the synthesizer to eliminate samples containing faults. For more details, please refer to Appendix A. We also provide an example of shared context in Appendix F.
3.4 Passage-level Utility Attribution
Specifically, the context $D$ recalled by the upstream retriever consists of $k$ passages. To evaluate the utility of each individual passage with inter-passage interactions, we adopt a perturbation-based method where we remove certain passages and inspect the changes in the final output. The approach is implemented via introducing a perturbation vector $\mathbf{v}\in\left\{0,1\right\}^{k}$, where 0 and 1 indicate whether the corresponding passage is removed or included, respectively. However, running all generations of $2^{k}$ possible perturbation vectors can result in significant computational overhead. Inspired by the method of Local Interpretable Model-agnostic Explanations (LIME; Ribeiro et al., 2016; Mardaoui and Garreau, 2021), we first sample $n$ perturbation vectors randomly and then fit a surrogate model for predicting the utility score, as shown below:
where we adopt the ridge regression (Hilt and Seegrist, 1977) as our surrogate model, $\bm{\alpha}$ represents the parameters to be fitted, $\lambda$ is a hyperparameter for regularization, and $z_{i}$ is the observed value under $\mathbf{v}_{i}$. More specifically, $\bm{\alpha}^{(i)}$ denotes the utility score of passage $d_{i}$, $\bm{\alpha}^{(0)}$ represents the intercept term. And $z_{i}$, which quantifies the fluctuation caused by $\mathbf{v}_{i}$, is calculated using the logit values of the tokens in the ground truth $\mathbf{y}$ at each time step, as shown below:
To evaluate the effectiveness of the above utility attribution method, we conduct a preliminary experiment on the GTI benchmark (Zhang et al., 2024), which includes three datasets: HotpotQA  (Yang et al., 2018), Natural Questions  (NQ; Kwiatkowski et al., 2019), and MSMARCO-QA  (Bajaj et al., 2018). Each test sample comprises input, ground truth, and ten passages including correct passages and other noise passages. We use the utility score to rank the passages. The results, measured using nDCG, demonstrate that our method shows a high accuracy in reflecting passage utility, as shown in Figure 2. We also compare our method to other attribution approaches, and our method outperforms them by over 20%. For further details of the experiment, please refer to Appendix B.
@@ 4.1 Datasets and Evaluation
We collect both in-domain and out-of-domain datasets for our experiments. In-domain datasets are utilized for providing seed data to construct synthetic training data, while out-of-domain datasets possess different tasks and corpora and are collected for further generalization tests. We collect seven datasets from KILT (Petroni et al., 2021), and three from BeIR (Thakur et al., 2021), as detailed in Table 1. All KILT datasets utilize Wikipedia dump dated 2019-08-01222http://dl.fbaipublicfiles.com/BLINK/enwiki-pages-articles.xml.bz2 as the corpus. Following Wang et al. (2019), we split the original articles into segments with a maximum length of 100 words, resulting in a total of 28,773,800 passages. For test sets of BeIR, we adopt their self-constructed corpora. For retrieval, we follow the closed corpus setup(Asai et al., 2023), where RALMs only retrieve from the corpus of the current dataset. For the test data, we randomly sample 1,000 data from the test split of each dataset.
For evaluation metrics, we mainly assess the performance of downstream tasks. For WoW, we use F1. For ELI5 and FiQA, we use ROUGE-L. For other datasets, we use accuracy.